OpenStudy (anonymous):

if the 7th term is in a geomteric sequence is 64 and the common ration is -2 what are the 1st term and the 14th term helpppp

OpenStudy (anonymous):

general term of a g s is a.r^(n-1) so a(-2)6 = 64 so a =?

OpenStudy (anonymous):

-12?

OpenStudy (anonymous):

sry wat i meant was (-2)^6

OpenStudy (anonymous):

so this would be the steps (-2)^6=64 and what would a =?

OpenStudy (anonymous):

thats the part that i dont get

OpenStudy (anonymous):

wat do u think?

OpenStudy (anonymous):

multiply -2 * 6 get the answere and then subtract it by 64?

OpenStudy (anonymous):

this a geometric prog.......so u dont subtract.........u just multiply nd divide stuff

OpenStudy (anonymous):

-5.333??? it cant be a decimal

OpenStudy (anonymous):

-.- ax64 = 64 so a=?

OpenStudy (anonymous):

768?

OpenStudy (anonymous):

how did u get that? a = 64/64 so a=?

OpenStudy (anonymous):

i multiplied 12 and 64

OpenStudy (anonymous):

the answere 1?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so for the 14 th term 1*(-2)^(14-1) = ?

OpenStudy (anonymous):

can u show the steps? please

OpenStudy (anonymous):

jus use that formula nth term in a gp is a*r*r*r*r....... (n-1) times

OpenStudy (anonymous):

-8192?

OpenStudy (anonymous):

is it that?

OpenStudy (anonymous):

how did u get -8192?

OpenStudy (anonymous):

it's (-2)^13

OpenStudy (anonymous):

yea u r right

OpenStudy (anonymous):

thank u so much ! and thank u for your time!